In which quadrants does the graph of the following inequality lie? 2x - y <= 4
a) I,II only
b) I,II,III only
c) I,III,IV only
d) I,II,III,IV
e) none of these

Answer by checkley71(8403)

2X-Y<=4
-Y=2X+4
Y=-2X-4 THEN THE Y INTERCEPT IS (0,-4)SETTING Y=0 WE FIND THE VALUE OF X
0=2X+4
2X=-4
X=-4/2
X=-2 OR (-2,0)
ANSWER D) NONE OF THE ABOVE.
THIS LINE FALLS WITH QUADRANTS II,III,& IV ONLY.